# ManiWAV: Learning Robot Manipulation from In-the-Wild Audio-Visual Data

### v1

@@ 4.2 Wiping Task
In this task, the robot is tasked to wipe a shape (e.g. heart, square) drawn on a whiteboard. The robot can start in any initial configuration above the whiteboard and grasp an eraser in parallel to the board. The main challenge of the task is that the robot needs to exert a reasonable amount of contact force on the whiteboard while moving the eraser along the shape. We collect 120 demonstrations and each demonstration takes about 10-15 seconds to complete.
@@
Test Scenarios: We run 20 rollouts for each policy. In addition to the T1 (5 / 20) and T2 (4 / 20) test cases as described above, we also test generalization to unseen table heights, erasers and drawing shape (11 / 20). A detailed breakdown of the test scenarios can be found in appendix.
@@
Fig 7: Pouring Evaluation. In the first row, we show the task definition and generalization scenarios to unseen number of dices and items. The second row shows the typical failure case for each baseline method and the overall task success rate. A breakdown of the substep success rate can be found in appendix.
@@ 4.3 Pouring Task
In this task, the robot is tasked to pick up the white cup and pour dice out to the pink cup if the white cup is not empty. When finish pouring, the robot needs to place the empty cup down to a designated location. The challenge of the task is that the robot cannot observe whether there are dice in the cup or not given the camera view point both before and after the pouring action. Therefore it needs to leverage a ‘shaking’ action to verify the cup is empty or not.
We collect 100 demonstrations for this task. It takes about 10-15 seconds to collect each demonstration. With the hand-held gripper, we can intuitively demonstrate a ‘shaking’ action that generates vibrations that can be captured by the contact microphone if there are dice inside the cup.
@@
Test Scenarios: We run 12 rollouts for each policy. In addition to the T1 and T2 scenarios, we also tested generalization scenarios (T3) with unseen number of dice (e.g., no dice or >>> 6 dices) and unseen objects (e.g., screws or beans).
Findings: The quantitative result and typical failure cases is visualized in Fig. 7. Our key findings in the experiments are: 1) When combined with information seeking action, audio can provide critical state information beyond visual observations. As illustrated in the figure, the vision only policy fails to execute the pour action as it cannot infer whether there’s dice in the cup or not, whereas the policy with audio feedback can leverage vibrations to infer this information and generalize to objects with similar sounds (e.g. screws). 2) Policy performance is sensitive to audio history length. As shown in the result, [1s audio] yields significantly lower performance since the shortened audio window does not contain sufficient information to guide robot actions. [10s audio] shows that even though longer audio window contains sufficient information, it increases the complexity of the learning process and require a higher capacity model.
Fig 8: Taping Evaluation. In the first row, we show the task definition. The second row shows the typical failure case for each method and the overall task success rate. A breakdown of the substep success rate can be found in appendix.
@@ 4.4 Taping Task
The robot is tasked to choose the ‘hook’ tape from several tapes (either ‘hook’ or ‘loop’) and strap wires by attaching the ‘hook’ tape to a ‘loop’ tape underneath the wires. We collect 200 demonstrations in total, with a ‘sliding’ primitive where we use the tip of the gripper finger to slide along the tape.
@@
  • Noise Reduction: Instead of applying training time noise augmentation, we evaluated an alternative method that uses a test-time noise reduction algorithm. The algorithm estimates a noise threshold for each frequency and applies a smoothed mask on the spectrogram [43]. More details can be found in appendix.
Test Scenarios: We run 10 rollouts in total, each time the robot is presented 2-4 velcro tapes in random order with at least one is a ‘hook’ tape.
Findings. 1) Contact microphone is sufficiently sensitive to different surface materials. As shown in Fig. 8, the vision only policy makes random decisions and yields a 20% success rate. Similarly the system uses environment microphone achieves similar result as the [Vision only] method, as it fails to pick up the subtle differences between the surface material. In contrast, by leveraging the contact microphone, our method is able to reliably guide the robot to pick up the correct tape. 2) Training-time noise augmentation is more effective than test-time noise reduction. This is because the noise cancellation algorithm causes the signal to be deprecated, resulting in domain gap between training and testing. On the other hand, our noise augmentation method preserves the frequency distribution of the original signals. Visualization can be found in appendix.
@@ 4.5 Limitations and Future Directions
On the data front, even though the contact microphone can pick up a wide range of audio signals, it may not be useful in certain scenarios (e.g. for deformable objects such as cloth or certain quasi-static tasks) where the interaction does not generate salient signals, and can easily become imperceptible due to robot motor noises during deployment. On the policy learning front, while we find the current policy network architecture achieves good performance on several tasks, it does not leverage the fact that audio signal are received at a higher frequency than images and can be used to learn more reactive behaviors. Future work can consider a hierarchical network architecture [44] that infers higher frequency actions from audio inputs.
@@ 5 Conclusion
Audio signals reveal rich information about the robot interaction and object properties, which can ease the learning of contact-rich robot manipulation policies. We present ManiWAV, an in-the-wild audio-visual data collection and policy learning framework. We design an ‘ear-in-hand’ gripper to collect human demonstrations that can be directly used to train robot policies through behavior cloning. By learning an effective audio-visual representation as condition to the diffusion policy, our method outperforms several alternative approaches on four contact-rich manipulation tasks, and generalizes to unseen in-the-wild environments.
@@ Acknowledgments
The authors would like to thank Yifan Hou and Zhenjia Xu for their help with discussions and setup of real-world experiments, Karan Singh for assistance with audio visualization and data collection. In addition, we would like to thank all REALab members: Huy Ha, Mandi Zhao, Mengda Xu, Xiaomeng Xu, Chuer Pan, Austin Patel, Yihuai Gao, Haochen Shi, Dominik Bauer, Samir Gadre for fruitful technical discussions and emotional support. The authors would also like to specially thank Xiaoran ‘Van’ Fan for his help with task brainstorm and audio expertise. This work was supported in part by the Toyota Research Institute, NSF Award #2143601, #2132519 and Sloan Fellowship. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ A.1 Audio Latency Calibration
Similar to using a clapperboard in film production, we tap on the contact microphone and match the time between the frame when the finger is observed to be in contact with the microphone and the corresponding audio signal captured by the contact microphone. A visual illustration is shown in Fig. 9. The audio is received about 0.06s after the image is received, as a result the total audio latency is 0.17 + 0.06 = 0.23s, where 0.17s is the calibrated image latency following the approach in [34].
@@ A.2 Model Details
Transformer Encoder. We use one transformer encoder layer with 8 heads to fuse the vision and audio features. We set feedforward dimension to 2048 and dropout ratio to $0.$
@@ B.2.1 Pouring Task
A breakdown of the success rate for each substep in the pouring task is shown in Tab. 1. ‘Grasp’ is successful if the robot grasps the white cup stably in its end effector, ‘Pour’ is successful if the robot pours all objects in the white cup to the pink cup on the table, ‘Place’ is successful if the robot places the white cup on the table after pouring. By using audio feedback to infer the object state (whether there’s object in the cup or not), our method is able to reliably guide the policy to pour objects and place the cup, whereas the baselines either never executes the pour action or the place action, resulting in low substep success rate. Videos of the policy rollouts can be found on the project website.
@@ B.2.2 Taping Task
A breakdown of the success rate for each substep in the taping task is shown in Tab. 2. ‘Touch’ is successful if the robot slides along the tape while maintaining contact, ‘Sense’ is successful if the robot chooses the correct tape, ‘Pick’ is successful if the robot successfully grasps the tape. ‘Place’ is successful if the robot successfully places the tape on top of the wires. By leveraging audio feedback to infer the object surface material (whether the tape is a ‘hook’ or ‘loop’), our method is able to reliably guide the policy to choose the correct tape whereas the baselines make random decisions, as shown in the ‘Sense’ step success rate. Videos of the policy rollouts can be found on the project website.
@@
Noise Reduction Algorithm. We use the non-stationary noise reduction method introduced in [43]. The algorithm computes a spectrogram from the audio waveform, and then apply IIR filter forward and backward on each frequency channel to obtain a time-smoothed version of the spectrogram. A mask is then computed base on the spectrogram from estimating a noise threshold for each frequency band of the signal/noise. And finally, a smoothed, inverted version of the mask is applied on the original spectrogram to cancel noise. In our experiments, we apply this algorithm directly on the real time audio signals before feeding it to the model for inference. In Fig. 11, we show spectrogram visualization of the real time signal (b) and the signal after reduction (c). Even though the noise reduction seems to successfully remove most of the robot noises and some other background noises, it does not preserve the original signal well, and still result in domain gap as comparing to the training signal (a).
@@ 19 T3 (-2 inch) 1 0 (float) 0 (incomplete) 0 (press) 1
‘Float’ means the robot keeps floating above the board without contacting the board before it wipes off the shape. ‘Press’ means the robot exerts too much force downward and causes the gripper to bend against the board. ‘Incomplete’ means the robot fails to follow the shape, either stops wiping early or wipes in wrong location.
As we can see from Tab. 3, ‘Float’ and ‘Press’ are most common in the [Vision only] baseline especially when the table height is different than training, likely due to the fact that it’s insufficient to infer contact from the top-down view wrist-mount camera image alone (as shown in Fig. 12).
@@
The most common failure case in [MLP fusion] is ‘incomplete’, where the robot stops wiping and releases the eraser before the shape is completely wiped off. We hypothesize that this is because simply fusing vision and audio features with MLP layers lose information that’s crucial for inferring the stage of the task.
@@ 19 T3 (unseen height) 1 0 (displace) 0 (lose) 0 (displace) 0 (displace)
‘Poke’ means that the robot pokes the spatula on the side of the bagel instead of inserting the spatula between the pan and the bottom of the bagel. ‘Lose’ means the robot loses contact with the bagel before it is flipped, as a result, the bagel falls back to its original side. ‘Displace’ means that the spatula is displaced in the robot end effector as comparing to its initial pose, as a result of the robot keeps moving down instead of switching to slide the spatula along the bottom of the pan.
For all T1 scenarios, we randomize the robot initial pose and object positions (e.g. bagel and pan). We can observe that the [Vision only] policy does not generalize well across initial configurations and most failure cases are either poking on the side of the bagel (since it’s hard to infer from the image alone if the spatula is contacting the bottom of the pan), or losing contact with the bagel early before it can be flipped.
Using a [ResNet] and [AVID] audio encoder results in the spatula to ‘displace’ most of the times, likely because the model is not sensitive enough to the sound feedback of spatula touching the bottom of the pan, and as a result keeps moving downward and causes the spatula to displace.

### v2

@@ 4.2 Wiping Task
In this task, the robot is tasked to wipe a shape (e.g., heart, square) drawn on a whiteboard. The robot can start in any initial configuration above the whiteboard and grasp an eraser in parallel to the board. The main challenge of the task is that the robot needs to exert a reasonable amount of contact force on the whiteboard while moving the eraser along the shape. We collect 119 demonstrations in total for this task.
@@
Test Scenarios: We run 20 rollouts for each policy. In addition to the T1 (5 / 20) and T2 (4 / 20) test cases as described above, we also test generalization to unseen table heights, erasers, and drawing shape T3 (11 / 20). A detailed breakdown of the test scenarios can be found in the appendix.
@@
Fig 7: Pouring Evaluation. In the first row, we show the task definition and generalization scenarios for an unseen number of dice and items. The second row shows the typical failure case for each baseline method and the overall task success rate. A breakdown of the substep success rate can be found in the appendix.
@@ 4.3 Pouring Task
The robot is tasked to pick up the white cup and pour dice out to the pink cup if the white cup is not empty. When finished pouring, the robot needs to place the empty cup down to a designated location. The challenge of the task is that the robot cannot observe whether there are dice in the cup or not given the camera viewpoint both before and after the pouring action. We collect 145 demonstrations for this task, with a ‘shaking’ action to generate vibrations that can be captured by the contact microphone if there are dice inside the cup.
@@
Test Scenarios: We run 12 rollouts for each policy. In addition to the T1 and T2 scenarios, we also tested generalization scenarios (T3) to an unseen number of dice (e.g., no dice or >>> 6 dice) and unseen objects (e.g., screws or beans).
Findings: The quantitative result and typical failure cases are visualized in Fig. 7. Our key findings in the experiments are: 1) Combining with information-seeking action, audio can provide critical state information beyond visual observations. As illustrated in the figure, the vision only policy fails to execute the pour action as it cannot infer whether there are dice in the cup or not, whereas the policy trained with audio feedback can leverage vibrations to infer the state information and generalize to objects with similar sounds (e.g., screws). 2) Policy performance is sensitive to audio history length. As shown in the result, [1s audio] yields significantly lower performance since the shortened audio window does not contain sufficient information to guide robot actions. [10s audio] shows that even though a longer audio window contains sufficient information, it increases the complexity of the learning process and requires a higher capacity model.
Fig 8: Taping Evaluation. In the first row, we show the task definition. The second row shows the typical failure case for each method and the overall task success rate. A breakdown of the substep success rate can be found in the appendix.
@@ 4.4 Taping Task
The robot is tasked to choose the ‘hook’ tape from several velcro tapes (either ’hook’ or ’loop’) and strap wires by attaching the ’hook’ tape to a ‘loop’ tape underneath the wires. We collect 193 demonstrations in total, with a ’sliding’ primitive where we use the tip of the gripper finger to slide along the tape. We run 10 rollouts in total, each time the robot is presented 2-4 velcro tapes in random order with at least one ’hook’ tape.
@@
  • Noise Reduction: Instead of applying training time noise augmentation, we evaluated an alternative method that uses a test-time noise reduction algorithm. The algorithm estimates a noise threshold for each frequency and applies a smoothed mask on the spectrogram [46]. More details can be found in the appendix.
Findings. 1) Contact microphone is sufficiently sensitive to different surface materials. As shown in Fig. 8, the [Vision only] policy makes random decisions and yields a 20% success rate. Similarly, the system that uses environment microphone to collect audio data achieves similar results as the [Vision only] method, as it fails to pick up the subtle differences between the surface material. In contrast, by leveraging the contact microphone, our method is able to reliably guide the robot to pick up the correct tape. 2) Training-time noise augmentation is more effective than test-time noise reduction. This is because the noise cancellation algorithm causes the signal to be deprecated, resulting in a domain gap between training and testing. On the other hand, our noise augmentation method preserves the frequency distribution of the original signals. More visualization can be found in the appendix.
@@ 4.5 Limitations and Future Directions
On the data front, even though the contact microphone can pick up a wide range of audio signals and is robust against environment noises in the background by design, it may not be useful in scenarios where the interaction does not generate salient signals (e.g., for deformable objects such as cloth or quasi-static tasks), and can easily become imperceptible due to robot motor noises during deployment. On the policy front, the current method does not leverage the fact that audio signals are received at a higher frequency than images and can be used to learn more reactive behaviors. Future work can consider a hierarchical network architecture [47] that infers higher frequency actions from audio inputs.
@@ 5 Conclusion
Audio signals reveal rich information about the robot interaction and object properties, which can ease the learning of contact-rich robot manipulation policies. We present ManiWAV, an in-the-wild audio-visual data collection and policy learning framework. We design an ‘ear-in-hand’ gripper to collect human demonstrations that can be directly used to train robot policies through behavior cloning. By learning an effective audio-visual representation as condition to the diffusion policy, our method outperforms several alternative approaches on four contact-rich manipulation tasks and generalizes to unseen in-the-wild environments.
@@ Acknowledgments
The authors would like to thank Yifan Hou and Zhenjia Xu for their help with discussions and setup of real-world experiments, Karan Singh for assistance with audio visualization and data collection. In addition, we would like to thank all REALab members: Huy Ha, Mandi Zhao, Mengda Xu, Xiaomeng Xu, Chuer Pan, Austin Patel, Yihuai Gao, Haochen Shi, Dominik Bauer, Samir Gadre for fruitful technical discussions and emotional support. The authors would also like to thank Xiaoran ‘Van’ Fan especially for his help with task brainstorming and audio expertise. This work was supported in part by the Toyota Research Institute, NSF Award #2143601, #2132519, and Sloan Fellowship. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.
@@ A.1 Audio Latency Calibration
Similar to using a clapperboard in film production, we tap on the contact microphone and match the time between the frame when the finger is observed to be in contact with the microphone and the corresponding audio signal captured by the contact microphone. A visual illustration is shown in Fig. 9. The audio is received about 0.06s after the image is received. As a result, the total audio latency is 0.17 + 0.06 = 0.23s, where 0.17s is the calibrated image latency following the approach in [1].
@@ A.2 Model Details
Spectrogram Parameters. There are several parameters that control the time and frequency resolution of the resulting spectrogram image when converting the audio waveform to a spectrogram. The window length is the length of the fixed intervals in which STFT divides the signal; it controls the time and frequency resolution of the spectrogram. Hop length is the length of the non-intersecting portion of the window lengths. The smaller the hop length is, the more times a particular audio segment will present in STFT, and the more elongated the time-axis of the resulting spectrogram will be. For a sample rate of 16kHz, we use the default window length, 400, as recommended by torchaudio. It is recommended that the hop length is around ½ of the window length; we choose 160 to slightly augment the signal pattern along the time-axis since contact signals are usually sparser in time compared to nature sounds or speech.
Transformer Encoder. We use one transformer encoder layer with 8 heads to fuse the vision and audio features. We set the feedforward dimension to 2048 and the dropout ratio to $0.$
@@ B.2.1 Pouring Task
A breakdown of the success rate for each substep in the pouring task is shown in Tab. 1. ‘Grasp’ is successful if the robot grasps the white cup stably in its end effector, ‘Pour’ is successful if the robot pours all objects in the white cup to the pink cup on the table, ‘Place’ is successful if the robot places the white cup on the table after pouring. By using audio feedback to infer the object state (whether there are objects in the cup or not), our method is able to reliably guide the policy to pour objects and place the cup, whereas the baselines either never execute the pour action or the place action, resulting in low substep success rate. We also find that the policy can generalize to unseen objects such as screws. In Fig. 11, we visualize the audio spectrogram when the robot ’shakes’ the cup. We can observe that the spectrogram for an empty cup is distinctive from a non-empty cup, and a cup with screws generates a similar audio pattern as a cup with dice upon shaking. We hypothesize that the audio features for the cup with screws and the cup with dice are also close in the audio feature space, leading to similar policy behavior. Videos of the policy rollouts can be found on the project website.
Fig 11: We visualize the audio spectrogram in the pouring task when the robot ’shakes’ the cup to sense whether there are objects inside the cup. (a) shows the audio feedback of an empty cup, (b) and (c) shows the audio feedback for a cup with dice and with screws inside, respectively.
@@ B.2.2 Taping Task
A breakdown of the success rate for each substep in the taping task is shown in Tab. 2. ‘Touch’ is successful if the robot slides along the tape while maintaining contact, ‘Sense’ is successful if the robot chooses the correct tape, ‘Pick’ is successful if the robot successfully grasps the tape. ‘Place’ is successful if the robot successfully places the tape on top of the wires. By leveraging audio feedback to infer the object surface material (whether the tape is a ‘hook’ or ‘loop’), our method is able to reliably guide the policy to choose the correct tape, whereas the baselines make random decisions, as shown in the ‘Sense’ step success rate. Videos of the policy rollouts can be found on the project website.
Noise Reduction Algorithm. We use the non-stationary noise reduction method introduced in [46]. The algorithm computes a spectrogram from the audio waveform, and then applies an IIR filter forward and backward on each frequency channel to obtain a time-smoothed version of the spectrogram. A mask is then computed based on the spectrogram by estimating a noise threshold for each frequency band of the signal/noise. Finally, a smoothed, inverted version of the mask is applied to the original spectrogram to cancel noise. In our experiments, we apply this algorithm directly to the real-time audio signals before feeding it to the model for inference. In Fig. 12, we show spectrogram visualization of the real-time signal (b) and the signal after reduction (c). Even though the noise reduction seems to successfully remove most of the robot noises and some other background noises, it does not preserve the original signal well and still results in a domain gap as compared to the training signal (a).
@@ 19 T3 (-2 inch) 1 0 (float) 0 (incomplete) 0 (press) 1
‘Float’ means the robot keeps floating above the board without contacting the board before it wipes off the shape. ‘Press’ means the robot exerts too much force downward and causes the gripper to bend against the board. ‘Incomplete’ means the robot fails to follow the shape, either stops wiping early or wipes in the wrong location.
As we can see from Tab. 3, ‘Float’ and ‘Press’ are most common in the [Vision only] baseline, especially when the table height is different than training, likely due to the fact that it’s insufficient to infer contact from the top-down view wrist-mount camera image alone (as shown in Fig. 13).
@@
The most common failure case in [MLP fusion] is ‘incomplete’, where the robot stops wiping and releases the eraser before the shape is completely wiped off. We hypothesize that this is because simply fusing vision and audio features with MLP layers loses information that’s crucial for inferring the stage of the task.
@@ 19 T3 (unseen height) 1 0 (displace) 0 (lose) 0 (displace) 0 (displace)
‘Poke’ means that the robot pokes the spatula on the side of the bagel instead of inserting the spatula between the pan and the bottom of the bagel. ‘Lose’ means the robot loses contact with the bagel before it is flipped. As a result, the bagel falls back to its original side. ‘Displace’ means that the spatula is displaced in the robot end effector as compared to its initial pose, as a result of the robot keeps moving down instead of switching to slide the spatula along the bottom of the pan.
For all T1 scenarios, we randomize the robot’s initial pose and object positions (e.g., bagel and pan). We can observe that the [Vision only] policy does not generalize well across initial configurations and most failure cases are either poking on the side of the bagel (since it’s hard to infer from the image alone if the spatula is contacting the bottom of the pan), or losing contact with the bagel early before it can be flipped.
Using a [ResNet] and [AVID] audio encoder results in the spatula to ‘displace’ most of the times, likely because the model is not sensitive enough to the sound feedback of spatula touching the bottom of the pan, and as a result the robot keeps moving downward and causes the spatula to displace.